OpenStudy (anonymous):

Solve: 6(x + 4) = 8(x − 2)

OpenStudy (anonymous):

first distribute then combine like terms

OpenStudy (homeworksucks):

As I said distribute so 6x+24=8x-16 and 16 to both sides so 6x+40=8x move the 6x over so 2x=40 x=20

OpenStudy (anonymous):

6x+24=8x-16 then combine

OpenStudy (anonymous):

8x-6x=2x, 24+16=40 x=20